# Variational formulations of transport on combinatorial meshes

### v6

@@ 5.1 Primal weak formulation
Formulation 5.4 (CMC transient primal weak formulation for transport phenomena).
@@
Formulation 5.5 (CMC steady-state primal weak formulation for transport phenomena).
Take the time-independent versions of assumptions in 5.2 and the operators in Equation 5.1. The unknown variable the potential $u[\mathsf{Y}]\in C^{0}\mathcal{K}$. We are solving the following system for $u$:
@@ 5.2 Mixed weak formulation
Formulation 5.6 (CMC transient mixed weak formulation for transport phenomena).
@@
$\forall r[\mathsf{X}\mathsf{T}^{-1}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},D-1},\;$ $A(r,q)-B_{\mathcal{D}}^{T}(r,\widetilde{u})-B_{\mathcal{A}}^{T}(r,\widetilde{u})$ $=-G(r)\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.7a)
$\forall\widetilde{w}[\mathsf{Y}\mathsf{L}^{D}]\in C^{D}\mathcal{K},\;$ $-B_{\mathcal{D}}(\widetilde{w},q)-C(\widetilde{w},\frac{\partial\widetilde{u}}{\partial t})$ $=-F(\widetilde{w})\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.7b)
@@
Formulation 5.7 (CMC steady-state mixed weak formulation for transport phenomena).
Take the time-independent versions of assumptions in 5.2 and the operators in Equation 5.6. The unknown variables are:
@@
$\forall r[\mathsf{X}\mathsf{T}^{-1}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},D-1},\;$ $A(r,q)-B_{\mathcal{D}}^{T}(r,\widetilde{u})-B_{\mathcal{A}}^{T}(r,\widetilde{u})$ $=-G(r)\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.9a)
$\forall\widetilde{w}[\mathsf{Y}\mathsf{L}^{D}]\in C^{D}\mathcal{K},\;$ $-B_{\mathcal{D}}(\widetilde{w},q)$ $=-F(\widetilde{w})\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.9b)
@@ 5.3 Comparison with other methods
Discussion 5.8.
Two methods that are notoriously based on differential geometry and meshes (i.e, forms and cochains) are discrete exterior calculus (DEC, [36]) and finite element exterior calculus (FEEC, [30]).
DEC is usually presented as a discretisation method based on simplicial meshes and their circumcentric dual, used for producing diagonal Hodge star matrices, where Hodge star operators act between primal $p$-cochains and dual $(D-p)$-cochains. The method does not define the full metric calculus (inner product, adjoint coboundary operator) and cannot be applied to our mixed formulations.
FEEC is a formalism for creating finite element spaces in a structured way based on smooth exterior calculus. It is purely a discretisation method and cannot accommodate material properties of microstructural features of different dimensions. Polynomial finite elements are generally constructed via polynomial mappings of reference elements and cannot handle general curvilinear meshes, although a large class of such elements can be handled with polynomial approximations [37]. Virtual finite elements, on the other hand, can handle general shapes [38].
@@ 6 Simulation examples
Consider a continuous steady-state problem (3.23 or 3.27) defined on a $D$-dimensional manifold $M$, with $\partial M=\Gamma_{\mathbf{D}}\cup\Gamma_{\mathbf{N}}$ (Equation 3.17), and input data $(\kappa,f,g_{\mathbf{D}},g_{\mathbf{N}})$ (time-independent versions from Table 2). Let $\mathcal{M}$ be a mesh that realises $M$ through an embedding $\varphi\colon\mathcal{M}\to\mathcal{P}(M)$, such that all of its boundary cells lie in full in $\Gamma_{\mathbf{D}}$ or $\Gamma_{\mathbf{N}}$, i.e., it is not possible for a cell to have non-empty intersection with both of $\Gamma_{\mathbf{D}}$ and $\Gamma_{\mathbf{N}}$, while not being on their common boundary. Take the Forman decomposition $\mathcal{K}$ of $\mathcal{M}$, and by $\Gamma_{\mathbf{D}}^{\mathcal{K}}$ and $\Gamma_{\mathbf{N}}^{\mathcal{K}}$ denote the discrete versions of $\Gamma_{\mathbf{D}}$ and $\Gamma_{\mathbf{N}}$ respectively. Assume that $\kappa$ is constant everywhere, so its discrete counterpart $\kappa^{\mathcal{K}}$ has the same value on all $(D-1)$-cells (or $\widetilde{\kappa}^{\mathcal{K}}$ on $1$-cells). $f$, $v$, $g_{\mathbf{D}},$ and $g_{\mathbf{N}}$ are discretised using the corresponding de Rham maps:
@@
Hence, we get the following discrete input for 5.5 and 5.7:
@@
In the examples, $v^{\mathcal{K}}$ is assumed to be 00, i.e., we will not include advective terms.
@@
If internal production rate is supplied as a density, i.e., as 00-form $\tilde{f}[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}]$, we define $f:=\star_{0}\tilde{f}$.
If the prescribed flow rate is given as a density (a 00-form) $\widetilde{g_{\mathbf{N}}}[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{1-D}]$ on $\Gamma_{\mathbf{N}}$, we define $g_{\mathbf{N}}:=\star_{\Gamma_{\mathbf{N}},0}\widetilde{g_{\mathbf{N}}}$.
@@
Because $[[\frac{\partial}{\partial x_{p}}]]=\mathsf{L}^{-1}$ (since $[[dx^{p}]]=\mathsf{L}$ and vector fields are dual to forms), $[[{\bf v}]]=\mathsf{T}^{-1}$. (This may sound counter-intuitive but note that the components $\{v_{p}\}_{p=1}^{D}$ of ${\bf v}$ have the dimension of speed.) Next, define the flat operator (one of the musical isomorphisms, the other one being its inverse, the sharp operator), $\flat\colon\mathcal{X}M\to\Omega^{1}M$, as follows: for any $X,Y\in\mathcal{X}M$,
@@
We can solve the resulting problem (primal: Equation 5.4, or mixed: Equation 5.9) with standard methods to get the potential $u^{\mathcal{K}}\in C^{0}\mathcal{K}$ and the flow rate $q^{\mathcal{K}}\in C^{D-1}\mathcal{K}$. Indeed, the operators and boundary conditions involved have the same forms for all discrete or discretised weak formulations, and so the LHS matrices, RHS vectors, and boundary constraints have the same form (see [39, Equations 2.7–2.10, 2.21–2.24]). However, one important feature of the mixed weak CMC formulation presented here is that the matrix $A$ is diagonal, and so the straightforward solution method of variable elimination works without introducing dense matrices, thus transforming a mixed LHS matrix to a a sparse symmetric positive definite one. Indeed, the mixed system is represented in the form $Aq-B^{T}u=-g,\ -Bq=-f$, and we can express $q$ from the first equation by $q=A^{-1}(-g+B^{T}u)$, which leads to
$-BA^{-1}(-g+B^{T}u)=-f\Rightarrow BA^{-1}B^{T}u=BA^{-1}g+f,$ (6.5)
@@
$u^{\mathcal{K}}$ is visualised trivially by applying it to the nodes of the mesh and creating a heat map. To visualise the $q^{\mathcal{K}}$ on a $(D-1)$-cell $S$ we proceed as follows. Let $S$ be boundary of $V_{-}$ and $V_{+}$ (if $S$ is on the boundary, we assume a ghost cell as the background). The direction of the flow is from $V_{+}$ to $V_{-}$ if $q^{\mathcal{K}}(S)>0$, from $V_{-}$ to $V_{+}$ if $q^{\mathcal{K}}(S)<0$, and there is no flow if $q^{\mathcal{K}}(S)=0$. The flow rate is then drawn as an arrow in the direction of the flow, with magnitude given by the heat map value of $q^{\mathcal{K}}(S)$ (no normalisation, e.g., dividing it by its measure $\mu_{D-1}(S)$, is done). No arrow is drawn if the flow is zero. We use a rainbow colour scheme, shown on Figure 1 (representing lowest value in red and highest value in magenta).
@@
    In Section 6.1 we will consider a $3$D example for a simple problem with exact potential on the unit cube. Application of the CMC formulations to more complex $3$D domains is left for future modelling of real microstructures.
@@
    In Section 6.4 we will simulate a problem with linear potential (and, hence, a constant flow rate) on a rectangle. We will, however test the CMC formulations on an irregular mesh generated by Neper ([40], https://neper.info).
@@ 6.2 Quadratic potential on a 2D disk
A regular polar mesh $\mathcal{M}$ for $M$ is constructed by dividing domains $[0,1]$ (of radius) and $[0,2\pi]$ (of rotation angle) into $4$ and $3$ equal parts respectively. The embedding of the Forman subdivision $\mathcal{K}$ is chosen so that the regularity of the mesh is preserved. $\mathcal{M}$ and $\mathcal{K}$ are shown on Figure 2. Exact solutions for the continuous problem on $M$ (discretised on $\mathcal{K})$ and solutions for the primal and mixed weak discrete formulations on $\mathcal{K}$ are visualised on Figure 3. The relative errors are: 0.0243588 for primal weak potential; 0.0802977 for mixed weak potential; 0.0581986 for primal weak flow rate; 4.72913e-06 for mixed weak flow rate.
@@ 6.3 Linear (in spherical coordinates) potential on a hemisphere
[…]
@@ 6.4 Linear potential on an irregular 2D mesh
In the first three examples we observed small relative errors for both potentials and flow rates. A common feature of those examples was that in the respective embedded meshes the topologically orthogonal sides were also geometrically orthogonal. This is not the case for the irregular mesh in the last example, suggesting that non-orthogonality of the embedded mesh is the reason for the high relative errors. Moreover, our tests on a problem with a parallelogram grid showed similarly high errors. A compelling reason for this is that the inner product on $1$-cells does not produce optimal Hodge star when topologically orthogonal edges are not geometrically orthogonal. This can be mended potentially by using non-diagonal inner products, i.e., by introducing additional measures for pairs of $p$-cells having a common boundary $(p-1)$-cell. Such an inner product will lead to non-local Hodge star (see [32, Section 8]) and adjoint coboundary operators but will still be computationally efficient. Indeed, the resulting linear systems for both primal and mixed formulations use only inner products, and can be expressed as sparse linear systems (although elimination of variables in the mixed formulation will not be effective). Regarding post-processing in the primal weak formulation, it can be reduced to solving a linear system with a sparse positive definite matrix, the matrix of the inner product.
@@ 7 Conclusions
In this article we developed a new formalism for modelling transport phenomena based on discrete cochains on quasi-cubical meshes that represent combinatorial differential forms on cell complexes with simple polytopes. We first described the exterior calculus formulation and using it as a benchmark formulated a discrete model where quantities and operators in exterior calculus are substituted with their discrete counterparts. However, it is clear that the derivation of the discrete formulation can be done without any reference to the continuum one and we believe this is a more natural approach. By using a computational mesh partition instead of the original mesh, we were able to: (a) express more complex cell interactions, and (b) use the already developed operators from [26] and [35]. The key benefit over the already explored applications [26, Section 4] is the use of variational primal and mixed formulations for the main model. The discrete variational formulations are very similar to those in the continuum for which finite (or virtual) elements are used, but in our case they are solved directly as the problems are already finite dimensional. We presented several numerical examples of discrete problems whose initial parameters are discretisation of continuum models, and demonstrated similarity to the discretised exact continuum solutions. Thus, although the primary aim of using CMC is for modelling micro-structures, we showed computationally (yet, without rigorous error analysis), that CMC formulations can also be viewed as new discretisation methods on quasi-cubical meshes with much less computational effort and with the ability to work directly on curved polytopes.
Some final remarks. A major challenge in materials modelling is the computational burden of solving governing equations at multiple scales. Cell complexes enable an efficient decomposition of space, allowing localized computations within individual cells while maintaining global consistency through their topological structure. This results in reduced computational overhead and improved parallelization capabilities. Moreover, the hierarchical nature of cell complexes aligns well with multigrid and multi-scale methods, making it easier to refine models adaptively based on physical requirements. The explicit topological structure of cell complexes allows for a more accurate representation of defects, such as dislocations, cracks, and voids. In continuum methods, discontinuities must be artificially introduced via enrichment techniques like extended finite element methods, while particle-based methods struggle with enforcing constraints. Cell complexes inherently represent discontinuities as boundaries between higher-dimensional elements, enabling more accurate and physically realistic modelling of defect dynamics. The work on formulation of conservation of momenta, required for modelling deformation and defect dynamics, is ongoing and will be reported in a future publication.

### v7

@@ 5.1 Primal weak formulation
Formulation 5.7 (CMC transient primal weak formulation for transport phenomena).
@@
Formulation 5.8 (CMC steady-state primal weak formulation for transport phenomena).
Take the time-independent versions of the assumptions in 5.2 and the operators in Equation 5.1. The unknown variable is the potential $u[\mathsf{Y}]\in C^{0}\mathcal{K}$. We are solving the following system for $u$:
@@ 5.2 Mixed weak formulation
Formulation 5.9 (CMC transient mixed weak formulation for transport phenomena).
@@
[…]
@@
Formulation 5.10 (CMC steady-state mixed weak formulation for transport phenomena).
Take the time-independent versions of the assumptions in 5.2 and the operators in Equation 5.6. The unknown variables are:
@@
[…]
@@ 5.3 Comparison with other methods
Discussion 5.11.
Two methods that are notoriously based on differential geometry and meshes (i.e, forms and cochains) are discrete exterior calculus (DEC, hirani2003discrete ) and finite element exterior calculus (FEEC, arnold2018finite ).
DEC is usually presented as a discretisation method based on simplicial meshes and their circumcentric dual, used for producing diagonal Hodge star matrices, where Hodge star operators act between primal $p$-cochains and dual (Dp)(D-p)-cochains. The method does not define the full metric calculus (inner product, adjoint coboundary operator) and cannot be applied to our mixed formulations.
FEEC is a formalism for creating finite element spaces in a structured way based on smooth exterior calculus. It is purely a discretisation method and cannot accommodate material properties of microstructural features of different dimensions. Polynomial finite elements are generally constructed via polynomial mappings of reference elements and cannot handle general curvilinear meshes, although a large class of such elements can be handled with polynomial approximations botti2018assesment . Virtual finite elements, on the other hand, can handle general shapes beirao2019virtual .
@@ 6 Simulation examples
Consider a continuous steady-state problem (3.23 or 3.27) defined on a $D$-dimensional manifold $M$, with $\partial M=\Gamma_{\mathbf{D}}\cup\Gamma_{\mathbf{N}}$ (Equation 3.17), and input data $(\kappa,f,g_{\mathbf{D}},g_{\mathbf{N}})$ (time-independent versions from Table 2). Consider a mesh $\mathcal{M}$ that realises $M$ through an embedding $\varphi\colon\mathcal{M}\to\mathcal{P}(M)$, such that all of its boundary cells lie in full in $\Gamma_{\mathbf{D}}$ or $\Gamma_{\mathbf{N}}$, i.e., it is not possible for a cell to have non-empty intersection with both of $\Gamma_{\mathbf{D}}$ and $\Gamma_{\mathbf{N}}$, while not being on their common boundary. Take the Forman decomposition $\mathcal{K}$ of $\mathcal{M}$, and by $\Gamma_{\mathbf{D}}^{\mathcal{K}}$ and $\Gamma_{\mathbf{N}}^{\mathcal{K}}$ denote the discrete versions of $\Gamma_{\mathbf{D}}$ and $\Gamma_{\mathbf{N}}$ respectively. Assume that $\kappa$ is constant everywhere, so its discrete counterpart $\kappa^{\mathcal{K}}$ has the same value on all (D1)(D-1)-cells (or $\widetilde{\kappa}^{\mathcal{K}}$ on $1$-cells). $f$, $v$, $g_{\mathbf{D}},$ and $g_{\mathbf{N}}$ are discretised using the corresponding de Rham maps:
@@
Hence, we get the following discrete input for 5.8 and 5.10:
@@
In the examples, $v^{\mathcal{K}}$ is assumed to be 0, i.e., we will not include advective terms.
@@
If internal production rate is supplied as a density, i.e., as 0-form $\tilde{f}[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}]$, we define $f:=\star_{0}\tilde{f}$.
If the prescribed flow rate is given as a density (a 0-form) $\widetilde{g_{\mathbf{N}}}[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{1-D}]$ on $\Gamma_{\mathbf{N}}$, we define $g_{\mathbf{N}}:=\star_{\Gamma_{\mathbf{N}},0}\widetilde{g_{\mathbf{N}}}$.
@@
Because $\frac{\partial}{\partial x_{p}}]]=\mathsf{L}^{-1}$ (since $dx^{p}]]=\mathsf{L}$ and vector fields are dual to forms), ${\bf v}]]=\mathsf{T}^{-1}$. (This may sound counter-intuitive but note that the components $\{v_{p}\}_{p=1}^{D}$ of $\bf v$} have the dimension of speed, $\mathsf{L}\mathsf{T}^{-1}$.) Next, define the flat operator (one of the musical isomorphisms, the other one being its inverse, the sharp operator), $\flat\colon\mathcal{X}M\to\Omega^{1}M$, as follows: for any $X,Y\in\mathcal{X}M$,
@@
We can solve the resulting problem (primal: Equation 5.4, or mixed: Equation 5.9) with standard methods to get the potential $u^{\mathcal{K}}\in C^{0}\mathcal{K}$ and the flow rate $q^{\mathcal{K}}\in C^{D-1}\mathcal{K}$. Indeed, the operators and boundary conditions involved have the same forms for all discrete or discretised weak formulations, and so the LHS matrices, RHS vectors, and boundary constraints have the same form (the details can be found in [berbatov2021guide, , Equations 2.7–2.10, 2.21–2.24]). However, one important feature of the mixed weak CMC formulation presented here is that the matrix $A$ is diagonal, and so the straightforward solution method of variable elimination works without introducing dense matrices, thus transforming a mixed LHS matrix to a a sparse symmetric positive definite one. Indeed, the mixed system is represented in the form $Aq-B^{T}u=-g,\ -Bq=-f$, and we can express $q$ from the first equation by $q=A^{-1}(-g+B^{T}u)$, which leads to
$BA^{-1}(-g+B^{T}u)=-f\Rightarrow BA^{-1}B^{T}u=BA^{-1}g+f,$ (6.5)
@@
$u^{\mathcal{K}}$ is visualised trivially by applying it to the nodes of the mesh and creating a heat map. To visualise the $q^{\mathcal{K}}$ on a (D1)(D-1)-cell $S$ we proceed as follows. Let $S$ be boundary of $V_{-}$ and $V_{+}$ (if $S$ is on the boundary, we assume a ghost cell as the background). The direction of the flow is from $V_{+}$ to $V_{-}$ if $q^{\mathcal{K}}(S)>0$, from $V_{-}$ to $V_{+}$ if $q^{\mathcal{K}}(S)<0$, and there is no flow if $q^{\mathcal{K}}(S)=0$. The flow rate is then drawn as an arrow in the direction of the flow, with magnitude given by the heat map value of $q^{\mathcal{K}}(S)$ (no normalisation, e.g., dividing it by its measure $\mu_{D-1}(S)$, is done). No arrow is drawn if the flow is zero. We use a rainbow colour scheme, shown on Figure 1 (representing lowest value in red and highest value in magenta).
@@
    In Section 6.1 we will consider a 33D example for a simple problem with exact potential on the unit cube. Application of the CMC formulations to more complex 33D domains is left for future modelling of real microstructures.
@@
    In Section 6.4 we will simulate a problem with linear potential (and, hence, a constant flow rate) on a rectangle. We will, however test the CMC formulations on an irregular mesh generated by Neper quey2011large , https://neper.info.
@@ 6.2 Quadratic potential on a 2D disk
A regular polar mesh $\mathcal{M}$ for $M$ is constructed by dividing domains $[0,1]$ (of radius) and $0,2\pi]$ (of rotation angle) into $4$ and $3$ equal parts respectively. The embedding of the Forman subdivision $\mathcal{K}$ is chosen so that the regularity of the mesh is preserved. $\mathcal{M}$ and $\mathcal{K}$ are shown on Figure 2. Exact solutions for the continuous problem on $M$ (discretised on $\mathcal{K})$ and solutions for the primal and mixed weak discrete formulations on $\mathcal{K}$ are visualised on Figure 3. The relative errors are: 0.0243588 for primal weak potential; 0.0802977 for mixed weak potential; 0.0581986 for primal weak flow rate; 4.72913e-06 for mixed weak flow rate.
@@ 6.3 Linear (in spherical coordinates) potential on a hemisphere
$(x,y,z)=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta),\ 0\leq\theta\leq\pi/2,\ 0\leq\varphi\leq 2\pi.$ (6.18)
@@ 6.4 Linear potential on an irregular 2D mesh
In the first three examples we observed small relative errors for both potentials and flow rates. A common feature of those examples was that in the respective embedded meshes the topologically orthogonal sides were also geometrically orthogonal. This is not the case for the irregular mesh in the last example, suggesting that non-orthogonality of the embedded mesh is the reason for the high relative errors. Moreover, our tests on a problem with a parallelogram grid showed similarly high errors. A compelling reason for this is that the inner product on $1$-cells does not produce optimal Hodge star when topologically orthogonal edges are not geometrically orthogonal. This can be mended potentially by using non-diagonal inner products, i.e., by introducing additional measures for pairs of $p$-cells having a common boundary (p1)(p-1)-cell. Such an inner product will lead to non-local Hodge star (see [wilson2007cochain, , Section 8]) and adjoint coboundary operators but will still be computationally efficient. Indeed, the resulting linear systems for both primal and mixed formulations use only inner products, and can be expressed as sparse linear systems (although elimination of variables in the mixed formulation will not be effective). Regarding post-processing in the primal weak formulation, it can be reduced to solving a linear system with a sparse positive definite matrix, the matrix of the inner product.
@@ 7 Conclusions
We have developed a new approach to modelling transport phenomena that bridges the gap between discrete microstructural topology and continuum physics. The Combinatorial Mesh Calculus (CMC) presented here provides the first variational formulations – both primal and mixed – for conservation laws on cell complexes, establishing a mathematical framework that is more faithful to material physics.
Theoretical significance: Our work demonstrates that transport phenomena can be formulated directly on discrete topological structures without reference to smooth manifolds. The exact preservation of conservation laws, discrete-continuous correspondences, and the natural treatment of multi-dimensional transport pathways represent a departure from approximating smooth equations to working intrinsically with the discrete topology of materials.
Computational impact: The diagonal structure of key matrices in the mixed formulation enables solution strategies unavailable to conventional methods. Our examples show that CMC handles curved geometries and irregular meshes naturally - features that challenge traditional finite elements. While the examples presented use manufactured solutions for verification, they demonstrate the method’s readiness for complex microstructural problems where different cell dimensions require distinct material properties.
Future directions: The CMC framework opens several research avenues:
  1. 1.
    Extension to momentum conservation for modelling deformation and defect dynamics (work in progress)
  2. 2.
    Coupling with experimental microstructural data through direct topology extraction
  3. 3.
    Multi-physics formulations leveraging the unified treatment of different conservation laws
  4. 4.
    Development of adaptive refinement strategies based on topological rather than geometric criteria
Broader implications: As materials science increasingly reveals the importance of microstructural features across scales, modelling approaches must evolve beyond the continuum-discrete dichotomy. CMC represents a mathematical framework aligned with physical reality – materials are neither collections of points nor smooth continua but complex assemblies of interacting components with distinct topological dimensions. By providing rigorous mathematics for this intermediate scale, we enable predictive modelling of phenomena from grain boundary engineering to defect-mediated transport. The computational efficiency demonstrated here, combined with the method’s intrinsic handling of complex geometries, positions CMC as a practical tool for materials design and analysis. The open source implementation ensures accessibility for both theoretical development and practical applications. This work establishes the foundation for a new generation of structure-aware computational methods in materials science.
CRediT authorship contrinution statement
Kiprian Berbatov: Writing – original draft, Writing – review & editing, Conceptualization, Methodology, Investigation, Software, Validation, Visualization. Andrey P. Jivkov: Writing – original draft, Writing – review & editing, Conceptualization, Methodology, Funding acquisition, Project administration.
Data availability
Code repository for this project can be found at https://github.com/kipiberbatov/cmc.
Declaration of competing interest
The authors declare no conflict of interests.
Acknowledgements
Jivkov acknowledges the financial support from the Engineering and Physical Sciences Research Council, UK, via grant EP/N026136/1.